Identify the critical t. An independent random sample is selected from an approximately normal population with unknown standard deviation. Find the degrees of freedom and the critical t value t∗t∗ for the given sample size and confidence level. Round critical t values to 4 decimal places.

Sample size, n Confidence level Degree of Freedom Critical value, t∗t∗
22 90
11 95
3 98
20 99
Answer #1

Solution: given thats df=n-1= 22-1 21 n 22 At g0 l Confidence level the tis de1-0.90 ty2,df- taos, 21 1.7027 df = At 95 Con fn 20 df -n-1 At 99 1. confidence levdl the tis 1 19 20-1- 1 /2=O.00S tipsdfto.o5, 19- 2.8609 Coitical Value t Degree Confiden